OpenStudy (anonymous):

Set up but do not solve for the appropriate particular solution yp for the differential equation y''+16y=3xsin(4x) using the Method of Undetermined Coefficients (primes indicate derivatives with respect to x). In your answer, give undetermined coefficients as A, B, etc.

OpenStudy (anonymous):

I also would like to know hot to solve this, I'm trying but can't get it.

OpenStudy (anonymous):

ive been trying for the last 2 hours and cannot get it

OpenStudy (anonymous):

for polynomial=Ax+B for trignometric =Csin(4x)+Dcos(4x) yp=(Ax+B)sin(4x)+(CX+D)cos(4x) is more appropriate guess

OpenStudy (anonymous):

well actually its been more than 2 hours! I got an answer but its wrong!

OpenStudy (anonymous):

Thats wrong :(

OpenStudy (anonymous):

@ranyai12 it is correct and works yp=(Ax+B)cos(4x)+(CX+D)sin(4x) 100% correct

OpenStudy (anonymous):

you can look at he following page from Paul online notes on Differential Equations

OpenStudy (anonymous):

But when I put it into the answer slot for the online hw it shows as incorrect

OpenStudy (anonymous):

i do not know why may be due to some error but m giving u surety it is correct .may be there answer is arranged somehow differently .

OpenStudy (anonymous):

idk that problem is horrendous! im going to ask my professor about it tomorrow! Thanks for your help though!

OpenStudy (anonymous):

I've also checked Paul online notes and missed that part, know I understand a little better that particular form of the right term,